OpenStudy (anonymous):

Solve the following system of equations. d + e = 6 d – e = 4 A. (3, –1) B. (5, 1) C. no solution D. infinite number of solutions

OpenStudy (anonymous):

d + e = 6 d - e = 4 ________ 2d = 10 d = 5 5 + e = 6 e = 1 (5, 1)